# Quantum annealing applications, challenges and limitations for optimisation problems compared to classical solvers

### v3

@@ 4 The unit commitment problem – An energy system application
With the release of the LeapHybridCQM solver [45], D-Wave opened itself up to a wider range of problems, such as Mixed Integer Linear Problems (MILP), with a wide spread of applications in various fields. To evaluate its performance, we benchmark D-Wave’s current performance on a real-world Unit Commitment problem, a complex optimisation challenge in the power sector that plays a crucial role in power system operation and planning. The Unit Commitment Problem involves determining the optimal scheduling of generating units—deciding when each unit should be turned on or off—over a given time horizon. The goal is to meet the forecasted load at minimum total production cost while satisfying a range of constraints, including power balance, reserve requirements, transmission limits, and individual unit constraints such as generation limits, minimum up/down times, and ramp rates [63]. For this purpose, a well-established MILP formulation of the Unit Commitment problem is chosen [64], which consists of both continuous and binary variables while maintaining a fully linear structure in its objective function and constraints. The problem is tested at three different scales as defined in Table 1. The objective function of a unit commitment model is as follows:
min $\text{min }\sum_{g\in G}\sum_{t\in T}\Bigg{(}\underbrace{c_{g}(t)+CP_{g}^{1}\,u_{g}(t)}_{i}+\underbrace{\sum_{s=1}^{S_{g}}\left(CS^{s}_{g}\delta^{s}_{g}(t)\right)}_{ii}\,\Bigg{)},$ (15)
@@
For the full-scale problem, the default pre-set minimum run time is set by D-Wave at 20s. In this limit, the hybrid solver was not able to find a single feasible solution, shown in Figure 6(a). We therefore solved the problem 6 times with an increased minimum solve time of 150s and 2 times of 400s, shown in Figure 6(b). For comparison, the near-optimal solution with a gap of 0.05% was found by Gurobi in 180s with a value of $1.23\times 10^{6}$. We can observe a clear increase in solution quality by the hybrid solver as in both cases, feasible solutions were found. The best feasible solution for the 150s case yielded, within 6 run,s an objective function value of $5.9\times 10^{6}$, while two runs with 400s run time yielded a value of $5.7\times 10^{6}$. This is a clear indication that the minimum run time chosen by the solver itself is not optimal and has to be adjusted specifically for the problem at hand, as a balance between computational time and solution quality. Next, an investigation was conducted to observe how D-Wave would perform if the problem was downsized, as stated in Table 1. First, the Reduced-1 problem with 12 time periods and 4 segments was examined, making it approximately one-fourth of the size of the initial problem. We kept the default minimum run-time setting of 10s, as this is already significantly longer than Gurobi’s solve time of 1.34s for the optimal solution without a gap. The solution distribution of 70 runs is shown in Figure 7(a), while the deviation of each run to the classical optimal solution is presented in Figure 7(b).
@@
Table 1: Parameters of the full scale and the two reduced versions of the Unit Commitment problem adjusted from [64] and using the $rts\_gmlc$ data set therein.
@@
Although D-Wave found at least one feasible solution in every run except for one, its best solution differs by approximately a factor of 10 from the optimal. The Reduced-2 problem is reduced even further to just 2 time periods and a single segment of linearisation. The solution distribution over 70 runs is shown in Figure 8(a) and the deviation from the optimal classical solution in Figure 8(b).
@@ 5 Conclusions
By increasing the complexity of the problem through systematically adding additional constraints, we conclude that the solution quality diverges more and more from the classical solvers for an increasing number of constraints at D-Wave’s scalability limit. Additionally, the increase in run time is so drastic that the hybrid solver shows no benefit at this scale. Although we did see some competitive computational times by D-Wave when handling quadratic constraints in Section 3.3, the hybrid solver was not able to find the optimal solution with a factor of approximately two off in most runs. It is for the BQP that D-Wave’s hybrid solver seems the most promising. Due to the mathematical structure of BQPs being so closely related to the Ising Hamiltonian of Equation 5, D-Wave’s hybrid solver performs better than all three classical solvers. The most competitive classical solver is IPOPT, but D-Wave does find slightly better solutions in a shorter amount of time. Compared with CPLEX, which due to its deterministic nature will find the optimal solution, D-Wave is superior in regards to computational time. Regarding the MILP unit commitment problem, the benchmark underscored the importance of fine-tuning the minimum run time to enhance the solution quality. Despite various adjustments and reduced problem cases, D-Wave’s LeapHybridCQMSolver was unable to surpass Gurobi in terms of computational time or solution quality. As this specific problem has no quadratic binary instance in either the objective function or in any of the constraints, the case study indicates that these instances are currently necessary to potentially show a computational benefit. While it is promising that D-Wave can directly solve full MILP problems using their hybrid solvers, the computational advantage appears to be limited to BQP problems.
@@ Acknowledgements
The $rts\_gmlc$ dataset used for the unit commitment problem can be found in [64]. All other datasets used and/or analyzed during the current study are available from the corresponding author upon reasonable request.

### v4

@@ 4 The unit commitment problem – An energy system application
With the release of the LeapHybridCQM solver [45], D-Wave opened itself up to a wider range of problems, such as Mixed Integer Linear Problems (MILP), with a wide spread of applications in various fields. To evaluate its performance, we benchmark D-Wave’s current performance on a real-world Unit Commitment problem, a complex optimisation challenge in the power sector that plays a crucial role in power system operation and planning. The Unit Commitment Problem involves determining the optimal scheduling of generating units—deciding when each unit should be turned on or off—over a given time horizon. The goal is to meet the forecasted load at minimum total production cost while satisfying a range of constraints, including power balance, reserve requirements, transmission limits, and individual unit constraints such as generation limits, minimum up/down times, and ramp rates [63]. For this purpose, a well-established MILP formulation of the Unit Commitment problem is chosen [64], which consists of both continuous and binary variables while maintaining a fully linear structure in its objective function and constraints. The problem is tested at three different scales as defined in Table˜1. The objective function of a unit commitment model is as follows:
$\text{min }\sum_{g\in G}\sum_{t\in T}\Bigg(\underbrace{c_{g}(t)+CP_{g}^{1}\,u_{g}(t)}_{i}+\underbrace{\sum_{s=1}^{S_{g}}\left(CS^{s}_{g}\delta^{s}_{g}(t)\right)}_{ii}\,\Bigg),$ (15)
@@
For the full-scale problem, the default pre-set minimum run time is set by D-Wave at 20s. In this limit, the hybrid solver was not able to find a single feasible solution, shown in Figure˜6(a). We therefore solved the problem 6 times with an increased minimum solve time of 150s and 2 times of 400s, shown in Figure˜6(b). For comparison, the near-optimal solution with a gap of 0.05% was found by Gurobi in 180s with a value of $1.23\times 10^{6}$. We can observe a clear increase in solution quality by the hybrid solver as in both cases, feasible solutions were found. The best feasible solution for the 150s case yielded, within 6 run,s an objective function value of $5.9\times 10^{6}$, while two runs with 400s run time yielded a value of $5.7\times 10^{6}$. This is a clear indication that the minimum run time chosen by the solver itself is not optimal and has to be adjusted specifically for the problem at hand, as a balance between computational time and solution quality. Next, an investigation was conducted to observe how D-Wave would perform if the problem was downsized, as stated in Table˜1. First, the Reduced-1 problem with 12 time periods and 4 segments was examined, making it approximately one-fourth of the size of the initial problem. We kept the default minimum run-time setting of 10s, as this is already significantly longer than Gurobi’s solve time of 1.34s for the optimal solution without a gap. The solution distribution of 70 runs is shown in Figure˜7(a), while the deviation of each run to the classical optimal solution is presented in Figure˜7(b).
@@
Table 1: Parameters of the full scale and the two reduced versions of the Unit Commitment problem adjusted from [64] and using the rts_gmlcrts\_gmlc data set therein.
@@
Although D-Wave found at least one feasible solution in every run except for one, its best solution differs by approximately a factor of 10 from the optimal. The Reduced-2 problem is reduced even further to just 2 time periods and a single segment of linearisation. The solution distribution over 70 runs is shown in Figure˜8(a) and the deviation from the optimal classical solution in Figure˜8(b).
@@ 5 Conclusions
By increasing the complexity of the problem through systematically adding additional constraints, we conclude that the solution quality diverges more and more from the classical solvers for an increasing number of constraints at D-Wave’s scalability limit. Additionally, the increase in run time is so drastic that the hybrid solver shows no benefit at this scale. Although we did see some competitive computational times by D-Wave when handling quadratic constraints in Section˜3.3, the hybrid solver was not able to find the optimal solution with a factor of approximately two off in most runs. It is for the BQP that D-Wave’s hybrid solver seems the most promising. Due to the mathematical structure of BQPs being so closely related to the Ising Hamiltonian of Equation˜5, D-Wave’s hybrid solver performs better than all three classical solvers. The most competitive classical solver is IPOPT, but D-Wave does find slightly better solutions in a shorter amount of time. Compared with CPLEX, which due to its deterministic nature will find the optimal solution, D-Wave is superior in regards to computational time. Regarding the MILP unit commitment problem, the benchmark underscored the importance of fine-tuning the minimum run time to enhance the solution quality. Despite various adjustments and reduced problem cases, D-Wave’s LeapHybridCQMSolver was unable to surpass Gurobi in terms of computational time or solution quality. As this specific problem has no quadratic binary instance in either the objective function or in any of the constraints, the case study indicates that these instances are currently necessary to potentially show a computational benefit. While it is promising that D-Wave can directly solve full MILP problems using their hybrid solvers, the computational advantage appears to be limited to BQP problems.
@@ Acknowledgements
The rts_gmlcrts\_gmlc dataset used for the unit commitment problem can be found in [64]. All other datasets used and/or analyzed during the current study are available from the corresponding author upon reasonable request.